What is the probability that a randomly assigned password (3 letter , non repeating)has only vowels in it ? Thanks
Answer by Alan3354(69443)

What is the probability that a randomly assigned password (3 letter , non repeating)has only vowels in it ?
------------------
If you mean in English, there are 5 vowels in the 26 letters.
---
The 1st is 5 of 26
Then 4 of 25
Then 3 of 24
----
5/26 * 4/25 * 3/24 = 1/26 * 4/5 * 1/8 = 1/26 * 1/5 * 1/2
= 1/260